# Benchmarking Offline Goal-Conditioned RL

### v1

@@ 3 How Have Prior Works Benchmarked Offline GCRL?
While many excellent offline GCRL algorithms have been proposed so far, the community currently lacks a standardized way to evaluate their performance, unlike other fields in RL (Brockman et al., 2016; Tassa et al., 2018; Fu et al., 2020; Terry et al., 2021). Tasks used by prior works in offline GCRL are often limited for providing a proper evaluation for various reasons. For example, many works directly use tasks in existing offline RL benchmarks (not necessarily designed for offline goal-conditioned RL), such as D4RL AntMaze and Kitchen (Fu et al., 2020). However, since the tasks are designed for single-task offline RL, they often evaluate their multi-task policies only on the single, original goal when using these tasks (Eysenbach et al., 2022; Park et al., 2023; Zheng et al., 2024b; Myers et al., 2024), which results in limited evaluation. Some employ online GCRL tasks provided by Plappert et al. (2018) (e.g., Fetch tasks) with policy-collected datasets (Ma et al., 2022; Yang et al., 2023), but these tasks are mostly atomic (e.g., single pick-and-place) and do not sufficiently address various challenges in offline GCRL, such as long-horizon reasoning and goal stitching. While Roboverse (Fang et al., 2022; Zheng et al., 2024a) provides pixel-based manipulation tasks, the tasks are still relatively atomic and it requires installing multiple, fragmented dependencies, making it less approachable for researchers. Some prior works construct bespoke tasks and datasets to individually study the important and specific features of their algorithms (e.g., stitching (Ghugare et al., 2024) and stochasticity (Myers et al., 2024)), but it is often not entirely clear how algorithms compare to one another or if the new capabilities of new methods (e.g., stitching) come at the loss of other capabilities. These limitations of previous evaluation tasks have motivated us to create a new benchmark. In this work, we introduce a set of diverse tasks that cover various challenges in offline GCRL, enabling a much more thorough and multi-faceted evaluation than previous tasks. We summarize the properties of the previous tasks and our new tasks in Table 1 and refer to Appendix B for further discussion on related work.
@@ 5 Challenges in Offline Goal-Conditioned RL
(2) Goal stitching: Another important challenge is to stitch the initial and final states of different trajectories to learn more diverse behaviors. We call this “goal stitching.” Goal stitching is different from “regular” stitching in offline RL, which applies only when the dataset is suboptimal. Unlike regular stitching, goal stitching applies even when the dataset only consists of optimal, expert trajectories, because we can often acquire more diverse goal-reaching behaviors by stitching multiple trajectories together, regardless of their optimality. For instance, an agent can stitch two atomic pick-and-place behaviors to sequentially move two objects in a single episode, even when the dataset does not contain any double pick-and-place behaviors. Goal stitching is crucial to learning diverse behaviors in many real-world applications with high behavioral diversity and large state spaces. In our benchmark, we introduce many tasks with large state spaces to assess the ability to perform goal stitching.
@@ 7 Environments, Tasks, and Datasets
We now introduce the environments, tasks, and datasets in our benchmark. They can be broadly categorized into three groups: locomotion, manipulation, and drawing. Most tasks support both state- and pixel-based observations, and each task accompanies five pre-defined state-goal pairs for evaluation. We also provide a separate validation dataset for each dataset. Videos are available at https://seohong.me/projects/ogbench.
@@ 7.1 Locomotion Tasks
In our benchmark, we substantially extend the original PointMaze and AntMaze tasks proposed by D4RL (Fu et al., 2020). Unlike the original D4RL tasks, which only support Point and Ant agents and do not challenge stitching [^4] See Ghugare et al. (2024) for the details about this point., stochasticity, or pixel-based control, we support Humanoid control, pixel-based observations, and multi-goal evaluation, while providing more challenging and diverse types of mazes and datasets. The supported maze types are as follows:
@@ 7.2 Manipulation Tasks
We provide a manipulation suite with three types of robotic manipulation tasks, Cube, Scene, and Puzzle, with diverse difficulties and complexities. They are designed to test the agent’s object manipulation, sequential generalization, and combinatorial generalization abilities. These environments are based on the MuJoCo simulator (Todorov et al., 2012) and a $6$-DoF UR5e robot arm (Zakka et al., 2022). On these tasks, we provide “play”-style datasets (play(Lynch et al., 2019) collected by non-Markovian expert policies with temporally correlated noise. To support more diverse types of research (e.g., dataset ablation studies), we additionally provide more noisy datasets (noisy) collected by Markovian expert policies with uncorrelated Gaussian noise, which we describe in Appendix A.
@@ 8.1 Algorithms
We benchmark six representative offline GCRL algorithms: goal-conditioned behavioral cloning (GCBC(Lynch et al., 2019; Ghosh et al., 2021), goal-conditioned implicit {V, Q}-learning (GCIVL and GCIQL(Kostrikov et al., 2022; Park et al., 2023), quasimetric RL (QRL(Wang et al., 2023), contrastive RL (CRL(Eysenbach et al., 2022), and hierarchical implicit Q-learning (HIQL(Park et al., 2023). GCBC is the simplest goal-conditioned behavioral cloning method. GCIVL and GCIQL are offline GCRL algorithms that approximates the optimal value function using an expectile regression (Newey & Powell, 1987). QRL is a non-traditional GCRL method that fits a quasimetric value function with a dual objective. CRL is a “one-step” RL algorithm that fits a Monte Carlo value function via contrastive learning and performs one-step policy improvement. HIQL is a hierarchical RL method that extracts a two-level hierarchical policy from a single GCIVL value function. For benchmarking, we perform a similar amount of hyperparameter tuning for each method to ensure fair comparison. We refer the reader to Appendices C and D for the details.
@@ 8.2 Benchmarking Results and Q&As
QRL $\mathbf{52}$ [^7] We note that Zheng et al. (2024b) use a different evaluation scheme based on the maximum performance over evaluation epochs. We report the average performance over the last three evaluation epochs (see Appendix D).  (Zheng et al., 2024b) $37$$\pm 14$ QRL $\mathbf{53}$7 (Zheng et al., 2024b) $38$$\pm 8$
@@
Q: Why can’t I just use D4RL AntMaze instead of OGBench AntMaze for offline GCRL?
@@
The results are quite remarkable: they show that having the right amount of noise (i.e., state coverage) is extremely important for achieving good performance. For example, the performance drops from $99\%$ to $6\%$ if there is no noise in expert actions, even on the most basic cube pick-and-place task. This suggests that, we may need to prioritize coverage much more than optimality when collecting datasets for offline GCRL in the real world as well, and failing to do so may lead to (surprising) failures in learning. Like action noise, we believe there are many other important properties of datasets that significantly affect performance. We believe that our fully transparent, controllable data-generation scripts can facilitate such scientific studies.
@@ 9 Research Opportunities
  • How should we train subgoal representations? Somewhat weirdly, HIQL struggles much more with state-based observations than pixel-based observations on manipulation tasks. We suspect this is related to subgoal representations, given that HIQL uses an additional learning signal from the policy loss to further train subgoal representations only in pixel-based environments (which we found does not help in state-based environments). HIQL uses a value function-based subgoal representation (Appendix C), but is there a better, more stable way to learn subgoal representations for hierarchical RL and planning?
@@
  • How can we combine expressive policies with GCRL? In Appendix A, we show that current offline GCRL methods often struggle with datasets collected by non-Markovian policies in manipulation environments. Handling non-Markovian trajectory data is indeed one of the major challenges in behavioral cloning, for which many recent BC-based methods have been proposed (Zhao et al., 2023; Chi et al., 2023). How can we incorporate these recent advancements in behavioral cloning into offline GCRL?
@@ Acknowledgments
We provide the full implementation details in Appendix D. We provide the code as well as the exact command-line flags to reproduce the entire benchmark table, datasets, and expert policies at https://github.com/seohongpark/ogbench.
Appendix A Additional Datasets
@@
Appendix B Prior Work in Goal-Conditioned RL
The problem of reaching any goal from any state has long been considered one of the central problems in reinforcement learning and sequential decision making (Kaelbling, 1993; Schaul et al., 2015; Andrychowicz et al., 2017), owing to its unsupervised nature, simplicity, and generality. There are many unique features of goal-conditioned RL that make it distinct from other (multi-task) RL problems, such as the presence of recursive subgoal structures, metric structures, and probabilistic interpretations. These intriguing properties have led to the development of diverse families of online and offline GCRL algorithms based on hindsight relabeling (Andrychowicz et al., 2017), hierarchical learning (Dayan & Hinton, 1992; Chane-Sane et al., 2021; Li et al., 2022; Park et al., 2023), planning (Savinov et al., 2018; Eysenbach et al., 2019; Nasiriany et al., 2019; Huang et al., 2019; Hoang et al., 2021; Kim et al., 2021), metric learning (Wang et al., 2023; Park et al., 2024b; Myers et al., 2024), dual optimization (Ma et al., 2022; 2023; Sikchi et al., 2024), weighted behavioral cloning (Yang et al., 2022; 2023; Hejna et al., 2023), generative modeling (Zeng et al., 2023; Hong et al., 2023), and contrastive learning (Eysenbach et al., 2022; Zheng et al., 2024a; b). In this work, we also consider the problem of offline goal-conditioned RL; however, instead of proposing a new algorithm, we introduce a new benchmark and reference implementations to facilitate and advance algorithms research in offline GCRL.
Appendix C Offline GCRL Algorithms

### v2

@@ 3 How Have Prior Works Benchmarked Offline GCRL?
While many excellent offline GCRL algorithms have been proposed so far, the community currently lacks a standardized way to evaluate their performance, unlike other fields in RL (Brockman et al., 2016; Tassa et al., 2018; Fu et al., 2020; Terry et al., 2021). Tasks used by prior works in offline GCRL are often limited for providing a proper evaluation for various reasons. For example, many works directly use tasks in existing offline RL benchmarks (not necessarily designed for offline goal-conditioned RL), such as D4RL AntMaze and Kitchen (Fu et al., 2020). However, since the tasks are designed for single-task offline RL, they often evaluate their multi-task policies only on the single, original goal when using these tasks (Eysenbach et al., 2022; Park et al., 2023; Zheng et al., 2024b; Myers et al., 2024), which results in limited evaluation. Some employ online GCRL tasks provided by Plappert et al. (2018) (e.g., Fetch tasks) with policy-collected datasets (Ma et al., 2022; Yang et al., 2023), but these tasks are mostly atomic (e.g., single pick-and-place) and do not sufficiently address various challenges in offline GCRL, such as long-horizon reasoning and goal stitching. While Roboverse (Fang et al., 2022; Zheng et al., 2024a) provides pixel-based manipulation tasks, the tasks are still relatively atomic and it requires installing multiple, fragmented dependencies, making it less approachable for researchers. Some prior works construct bespoke tasks and datasets to individually study the important and specific features of their algorithms (e.g., stitching (Ma et al., 2022; Ghugare et al., 2024; Wang et al., 2024) and stochasticity (Myers et al., 2024)), but it is often not entirely clear how algorithms compare to one another or if the new capabilities of new methods (e.g., stitching) come at the loss of other capabilities. These limitations of previous evaluation tasks have motivated us to create a new benchmark. In this work, we introduce a set of diverse tasks that cover various challenges in offline GCRL, enabling a much more thorough and multi-faceted evaluation than previous tasks. We summarize the properties of the previous tasks and our new tasks in Table 1 and refer to Appendix C for further discussion on related work.
@@ 5 Challenges in Offline Goal-Conditioned RL
(2) Goal stitching: Another important challenge is to stitch the initial and final states of different trajectories to learn more diverse behaviors. We call this “goal stitching.” Goal stitching is different from “regular” stitching in offline RL, which applies only when the dataset is suboptimal. Unlike regular stitching, goal stitching applies even when the dataset only consists of optimal, expert trajectories, because we can often acquire more diverse goal-reaching behaviors by stitching multiple trajectories together, regardless of their optimality. For instance, an agent can stitch two atomic pick-and-place behaviors to sequentially move two objects in a single episode, even when the dataset does not contain any double pick-and-place behaviors. Goal stitching is crucial to learning diverse behaviors in many real-world applications with high behavioral diversity and large state spaces. In our benchmark, we introduce many tasks with large state spaces to assess the ability to stitch goals.
@@ 7 Environments, Tasks, and Datasets
We now introduce the environments, tasks, and datasets in our benchmark. They can be broadly categorized into three groups: locomotion, manipulation, and drawing. We provide a separate validation dataset for each dataset, and most tasks support both state- and pixel-based observations. Videos are available at https://seohong.me/projects/ogbench.
Evaluation. Each task in OGBench accompanies five pre-defined state-goal pairs for evaluation (Appendix F). Performance is measured by the average success rate across the five evaluation goals. For each pre-define state-goal pair, we perform multiple rollouts with slightly randomized initial and goal states. In each evaluation episode, a goal $g\in{\mathcal{S}}$ (which is simply another state; see Section 2) is given to the agent, and the episode immediately terminates when the agent reaches the goal. Each task has its own goal success criterion, which we describe in Section E.1.
Variants. While OGBench is mainly designed for (unsupervised) offline goal-conditioned RL, it supports several other problem settings as well. Before introducing our environments and datasets, we briefly mention two variants of OGBench tasks for standard offline RL and supervised offline goal-conditioned RL.
  • singletask: Single-task variants are designed for standard (i.e., non-goal-conditioned) offline RL. To convert a goal-conditioned task into a standard reward-maximizing task, we fix an evaluation goal and relabel the dataset with the corresponding reward function. Each environment provides five single-task tasks (for each of the five evaluation goals), which brings the total number of single-task tasks to $\mathbf{410}$. See Section E.2 for the details.
  • oraclerep: Oracle representation variants are designed for “supervised” offline goal-conditioned RL. In this setting, we provide low-dimensional oracle goal representations that contain only relevant information for fulfilling the goal success criterion (e.g., only the $x$-$y$ position of the agent in maze navigation environments). This reduces the burden of goal representation learning and may potentially be useful for analysis purposes. See Section E.3 for details.
@@ 7.1 Locomotion Tasks
In our benchmark, we substantially extend the original PointMaze and AntMaze tasks proposed by D4RL (Fu et al., 2020). Unlike the original D4RL tasks, which only support Point and Ant agents and do not challenge stitching [^4] See Ghugare et al. (2024) for the details about this point.
, stochasticity, or pixel-based control, we support Humanoid control, pixel-based observations, and multi-goal evaluation, while providing more challenging and diverse types of mazes and datasets. The supported maze types are as follows:
@@ 7.2 Manipulation Tasks
We provide a manipulation suite with three types of robotic manipulation tasks, Cube, Scene, and Puzzle, with diverse difficulties and complexities. They are designed to test the agent’s object manipulation, sequential generalization, and combinatorial generalization abilities. These environments are based on the MuJoCo simulator (Todorov et al., 2012) and a $6$-DoF UR5e robot arm (Zakka et al., 2022). On these tasks, we provide “play”-style datasets (play(Lynch et al., 2019) collected by non-Markovian expert policies with temporally correlated noise. To support more diverse types of research (e.g., dataset ablation studies), we additionally provide more noisy datasets (noisy) collected by Markovian expert policies with uncorrelated Gaussian noise, which we describe in Appendix B.
@@ 8.1 Algorithms
We benchmark six representative offline GCRL algorithms: goal-conditioned behavioral cloning (GCBC(Lynch et al., 2019; Ghosh et al., 2021), goal-conditioned implicit {V, Q}-learning (GCIVL and GCIQL(Kostrikov et al., 2022; Park et al., 2023), quasimetric RL (QRL(Wang et al., 2023), contrastive RL (CRL(Eysenbach et al., 2022), and hierarchical implicit Q-learning (HIQL(Park et al., 2023). GCBC is the simplest goal-conditioned behavioral cloning method. GCIVL and GCIQL are offline GCRL algorithms that approximates the optimal value function using an expectile regression (Newey & Powell, 1987). QRL is a non-traditional GCRL method that fits a quasimetric value function with a dual objective. CRL is a “one-step” RL algorithm that fits a Monte Carlo value function via contrastive learning and performs one-step policy improvement. HIQL is a hierarchical RL method that extracts a two-level hierarchical policy from a single GCIVL value function. For benchmarking, we perform a similar amount of hyperparameter tuning for each method to ensure fair comparison. We refer the reader to Appendices D and E for the details.
@@ 8.2 Benchmarking Results and Q&As
QRL $\mathbf{52}$ [^7] We note that Zheng et al. (2024b) use a different evaluation scheme based on the maximum performance over evaluation epochs. We report the average performance over the last three evaluation epochs (see Appendix E).  (Zheng et al., 2024b) $37$$\pm 14$ QRL $\mathbf{53}$7 (Zheng et al., 2024b) $38$$\pm 8$
@@
Q: Why should I use OGBench AntMaze instead of the D4RL one?
@@
The results are quite remarkable: they show that having the right amount of noise (i.e., state coverage) is very important for achieving good performance. For example, the performance drops from $99\%$ to $6\%$ if there is no noise in expert actions, even on the most basic cube pick-and-place task. This suggests that, we may need to prioritize coverage much more than optimality when collecting datasets for offline GCRL in the real world as well, and failing to do so may lead to (surprising) failures in learning. Like action noise, we believe there are many other important properties of datasets that significantly affect performance. We believe that our fully transparent, controllable data-generation scripts can facilitate such scientific studies.
@@ 9 Research Opportunities
  • How should we train subgoal representations? Somewhat weirdly, HIQL struggles much more with state-based observations than pixel-based observations on manipulation tasks. We suspect this is related to subgoal representations, given that HIQL uses an additional learning signal from the policy loss to further train subgoal representations only in pixel-based environments (which we found does not help in state-based environments). HIQL uses a value function-based subgoal representation (Appendix D), but is there a better, more stable way to learn subgoal representations for hierarchical RL and planning?
@@
  • How can we combine expressive policies with GCRL? In Appendix B, we show that current offline GCRL methods often struggle with datasets collected by non-Markovian policies in manipulation environments. Handling non-Markovian trajectory data is indeed one of the major challenges in behavioral cloning, for which many recent BC-based methods have been proposed (Zhao et al., 2023; Chi et al., 2023). How can we incorporate these recent advancements in behavioral cloning into offline GCRL?
@@ Acknowledgments
We provide the full implementation details in Appendix E. We provide the code as well as the exact command-line flags to reproduce the entire benchmark table, datasets, and expert policies at https://github.com/seohongpark/ogbench.
Appendix A Limitations
While OGBench covers a number of challenges in offline goal-conditioned RL, such as long-horizon reasoning, goal stitching, and stochastic control, there exist other challenges that our benchmark does not address. For example, all OGBench tasks assume that the environment dynamics remain the same between the training and evaluation environments. Also, although several OGBench tasks (e.g., Cube, Puzzle, and Powderworld) require unseen goal generalization to some degree, our tasks do not specifically test visual generalization to entirely new objects. Finally, we have made several trade-offs to reduce computational cost and to focus the benchmark on algorithms research at the expense of sacrificing realism to some degree (e.g., the use of the transparent arm in manipulation environments, the use of synthetic (yet fully controllable) datasets, etc.). Nonetheless, we believe OGBench can spur the development of performant offline GCRL algorithms, which can then help researchers develop scalable data-driven unsupervised RL pre-training methods for real-world tasks.
[…]
Environment Dataset Type Dataset GCBC GCIVL GCIQL QRL CRL HIQL
cube-double-play-v0 $1$$\pm 1$ $36$$\pm 3$ $\mathbf{40}$$\pm 5$ $1$$\pm 0$ $10$$\pm 2$ $6$$\pm 2$
cube-triple-play-v0 $1$$\pm 1$ $1$$\pm 0$ $3$$\pm 1$ 00 ±$\pm 0$ $\mathbf{4}$$\pm 1$ $3$$\pm 1$
noisy cube-single-noisy-v0 $8$$\pm 3$ $71$$\pm 9$ $\mathbf{99}$$\pm 1$ $25$$\pm 6$ $38$$\pm 2$ $41$$\pm 6$
cube-double-noisy-v0 $1$$\pm 1$ $14$$\pm 3$ $\mathbf{23}$$\pm 3$ $3$$\pm 1$ $2$$\pm 1$ $2$$\pm 1$
cube-triple-noisy-v0 $1$$\pm 1$ $\mathbf{9}$$\pm 1$ $2$$\pm 1$ $1$$\pm 0$ $3$$\pm 1$ $2$$\pm 1$
cube-quadruple-noisy-v0 $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$
noisy scene-noisy-v0 $1$$\pm 1$ $\mathbf{26}$$\pm 5$ $\mathbf{26}$$\pm 2$ $9$$\pm 2$ $1$$\pm 1$ $\mathbf{25}$$\pm 4$
noisy puzzle-3x3-noisy-v0 $1$$\pm 0$ $42$$\pm 19$ $\mathbf{94}$$\pm 3$ 00 ±$\pm 0$ $30$$\pm 6$ $51$$\pm 11$
puzzle-4x4-noisy-v0 00 ±$\pm 0$ $20$$\pm 3$ $\mathbf{29}$$\pm 7$ 00 ±$\pm 0$ 00 ±$\pm 0$ $16$$\pm 4$
puzzle-4x5-noisy-v0 00 ±$\pm 0$ $\mathbf{19}$$\pm 0$ $\mathbf{19}$$\pm 0$ 00 ±$\pm 0$ $3$$\pm 2$ $5$$\pm 1$
puzzle-4x6-noisy-v0 00 ±$\pm 0$ $17$$\pm 2$ $\mathbf{18}$$\pm 2$ 00 ±$\pm 0$ $6$$\pm 3$ $2$$\pm 1$
noisy visual-cube-single-noisy-v0 $14$$\pm 3$ $75$$\pm 3$ $48$$\pm 3$ $10$$\pm 5$ $39$$\pm 30$ $\mathbf{99}$$\pm 0$
visual-cube-double-noisy-v0 $5$$\pm 1$ $17$$\pm 4$ $22$$\pm 2$ $6$$\pm 2$ $6$$\pm 3$ $\mathbf{59}$$\pm 3$
visual-cube-triple-noisy-v0 $16$$\pm 1$ $18$$\pm 1$ $12$$\pm 1$ $9$$\pm 4$ $16$$\pm 1$ $\mathbf{23}$$\pm 2$
visual-cube-quadruple-noisy-v0 $9$$\pm 0$ 00 ±$\pm 0$ $2$$\pm 2$ 00 ±$\pm 0$ $8$$\pm 2$ $\mathbf{12}$$\pm 8$
noisy visual-scene-noisy-v0 $13$$\pm 2$ $23$$\pm 2$ $12$$\pm 4$ $2$$\pm 0$ $15$$\pm 2$ $\mathbf{50}$$\pm 1$
noisy visual-puzzle-3x3-noisy-v0 $1$$\pm 1$ $20$$\pm 0$ $26$$\pm 4$ 00 ±$\pm 0$ $1$$\pm 1$ $\mathbf{70}$$\pm 6$
visual-puzzle-4x4-noisy-v0 $7$$\pm 3$ $47$$\pm 3$ $49$$\pm 7$ 00 ±$\pm 0$ $6$$\pm 2$ $\mathbf{84}$$\pm 4$
visual-puzzle-4x5-noisy-v0 $6$$\pm 1$ $14$$\pm 10$ $\mathbf{19}$$\pm 0$ 00 ±$\pm 0$ $7$$\pm 1$ $14$$\pm 10$
visual-puzzle-4x6-noisy-v0 $2$$\pm 1$ $12$$\pm 8$ $\mathbf{17}$$\pm 1$ 00 ±$\pm 0$ $2$$\pm 1$ $14$$\pm 2$
Appendix B Additional Datasets
@@
Appendix C Prior Work in Goal-Conditioned RL
The problem of reaching any goal from any state has long been considered one of the central problems in reinforcement learning and sequential decision making (Kaelbling, 1993; Schaul et al., 2015; Andrychowicz et al., 2017), owing to its unsupervised nature, simplicity, and generality. There are many unique features of goal-conditioned RL that make it distinct from other (multi-task) RL problems, such as the presence of recursive subgoal structures, metric structures, and probabilistic interpretations. These intriguing properties have led to the development of diverse families of online and offline GCRL algorithms based on hindsight relabeling (Andrychowicz et al., 2017), hierarchical learning (Dayan & Hinton, 1992; Chane-Sane et al., 2021; Li et al., 2022; Park et al., 2023), planning (Savinov et al., 2018; Eysenbach et al., 2019; Nasiriany et al., 2019; Huang et al., 2019; Hoang et al., 2021; Kim et al., 2021; Wang et al., 2024), metric learning (Wang et al., 2023; Park et al., 2024b; Myers et al., 2024), dual optimization (Ma et al., 2022; 2023; Sikchi et al., 2024), weighted behavioral cloning (Yang et al., 2022; 2023; Hejna et al., 2023), generative modeling (Zeng et al., 2023; Hong et al., 2023), and contrastive learning (Eysenbach et al., 2022; Zheng et al., 2024a; b). In this work, we also consider the problem of offline goal-conditioned RL; however, instead of proposing a new algorithm, we introduce a new benchmark and reference implementations to facilitate and advance algorithms research in offline GCRL.
Appendix D Offline GCRL Algorithms